# Cristiá, Capozucca and Rossi

### v1

@@ 2.4 Restricted quantifiers
We have already introduced restricted quantifiers (RQ) in $\{log\}$. We have seen foreach (restricted universal quantifier, RUQ) and exists (restricted existential quantifier, REQ), in the previous sections. Here we discuss them a little more deeply. First of all, RQ in $\{log\}$ are at least as expressive as set relation algebra (Cristiá and Rossi, 2024d, , Sect. 4.3), in part, due to the following characteristics. First, RQ admit quantification terms rather than simply variables. For example:
@@ 2.5 Types
Rooted in Prolog, $\{log\}$ provides essentially an untyped language based on untyped FOL and untyped set theory. In this way, a set such as $\{1,a,(3,b),\{x,q\}\}$ is perfectly legal in $\{log\}$. Untyped formalisms are not bad in themselves but types help to avoid some classes of errors (Lamport and Paulson,, 1999). For this reason, we recently defined a type system and implemented a typechecker as a $\{log\}$ component (Cristiá and Rossi, 2024b, ).
@@ 3 A logic language for state machines
As we have explained in the introduction, we have defined a language on top of $\{log\}$ for the description of state machines. The language is inspired in the B and Z specification languages. This language constrains programmers to use $\{log\}$ in a more restricted way but increasing the chances of automatic verification. We will introduce the main elements of the language by means of a running example. Some elements are deliberately left out of the paper for brevity—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.1).
@@
A $\{log\}$ state machine is a collection of declarations and predicates. Predicates and declarations are similar to Prolog’s. State machines transition between states defined by means of state variables. The first declaration is used to declare the state variables of the state machine. In the birthday book we have the following declaration:
@@
invariant(birthdayFun).birthdayFun(Birthday) :- pfun(Birthday).As can be seen, each invariant is preceded by an invariant declaration. The body of each invariant is a $\{log\}$ formula depending on the state variables. An invariant declaration is a declaration of intent. It remains to be proved that these predicates are indeed preserved by the operations of the specification—see Section 5.
@@
[…]
@@
addBirthdayOk(Known,Birthday,Name,Date,Known_,Birthday_) :-  Name nin Known &  un(Known,{Name},Known_) &  un(Birthday,{[Name,Date]},Birthday_).It is easy to see that the specification is given by providing pre and postconditions. For instance, the constraint Name nin Known is a precondition whereas un(Birthday,{[Name,Date]},Birthday_) is a postcondition stating that the value of Birthday in the next state is equal to the union of its value in the before state with {[Name,Date]}.
@@
operation(remind).remind(Birthday,Today,Names) :-  rres(Birthday,{Today},M) & dom(M,Names).where rres(R,B,S) is a constraint called range restriction whose interpretation is $S=\{(x,y)\in R\mid y\in B\}$777rres/3 is one of the relational operators that can be defined in terms of the classic set operators and the basic operators of set relation algebra.. Note that state variable Known is not an argument simply because the operation does not need it. Here Today is (supposed to be) an input whereas Names is (supposed to be) an output. This distinction is enforced by the user when the specification is executed. remind does not transition to a new state; it just outputs information. This is specified by not including the next-state variables in the head of the clause.
@@
3.1 Other elements of the language
The state machine specification language also supports parameters, axioms and user-defined theorems (Rossi and Cristiá,, 2025, Section 13.1.1). Parameters play the role of machine parameters and constants in B specifications and the role of variables declared in axiomatic definitions in Z. That is, parameters serve to declare the existence of some (global) values accessible to invariants and operations, but they cannot be changed by operations—i.e., there is no next-state value for a parameter. Axioms are used to state properties of parameters. User-defined theorems are used to state properties that can be deduced from axioms, invariants, operations or theorems that have already been declared.
@@ 3.2 Typing state machines
operation(remind).dec_p_type(rel(name,date),date,set(name))).    % type declarationremind(Birthday,Today,Names) :-  rres(Birthday,{Today},M) & dom(M,Names) & dec(M,rel(name,date)).The dec_p_type declaration provides the type for remind. The term remind(rel(name,date),date,set(name)) declares the type of each argument of the operation. For instance, rel(name,date) is the type of Birthday, date is the type of Today, and so on. The type of M is given by an explicit type declaration, dec(M,rel(name,date)). In turn, name and date are basic or uninterpreted types; rel(name,date) is the type of all binary relations with domain in name and range in date; and set(name) is the type of all sets whose elements are of type name. If t is a basic type then its elements are of the form t:$\langle elem\rangle$, where $elem$ is any Prolog atom or natural number.
@@ 4.1 The Next environment
Next is a component tightly integrated with $\{log\}$ that simplifies the execution and analysis of deterministic, fully instantiated functional scenarios. Although Next is less general than the approach shown in the next section, it considerably ease the job of users. For example, we can add a person’s name and birth date when the system is in its initial state, as follows:
{log}=> initial >> addBirthday(Name:alice,Date:may24).  Known = {alice}, Birthday = {[alice,may24]}initial refers to the predicate declared as such in the specification; >> (read then) imposes a sequential order in the execution. In this way, $\{log\}$ automatically fetches state variables when calling an operation and chains the after-state of an operation with the before-state of the next one. Users need to indicate values only for the input variables (e.g., Name:alice). With Next is easy to add two persons to the system:
{log}=> initial >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]).  Known = {alice,bob},  Birthday = {[alice,may24],[bob,nov05]}
The execution trace of the scenario can be analyzed by enclosing the initial state between square brackets:
{log}=> [initial] >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]).  Known = {},  Birthday = {}    ----> addBirthday(Name:alice,Date:may24)  Known = {alice},  Birthday = {[alice,may24]}    ----> addBirthday(Name:bob,Date:nov05)  Known = {alice,bob},  Birthday = {[alice,may24],[bob,nov05]}
More operations can be called by using >>:
{log}=> [initial] >>        addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]) >>        remind(Today:may24,Names).
@@ 4.2 Symbolic executions
$\{log\}$ can perform symbolic executions too (King,, 1976). That is, we can execute the prototype by providing some variables instead of constants as inputs—or a mixture of both. In this way we can draw more general conclusions about the specification. In this case, though, we cannot use Next because it requires variables involved in the scenario to be completely instantiated. Actually, in symbolic executions we need to analyze the constraints returned by $\{log\}$. The following scenario analyses how remind responds to a call where Birthday has at least two equal dates in its range:
{log}=> pfun({[X,Y],[Z,Y]/B}) & remind({[X,Y],[Z,Y]/B},Y,Names).  Names = {X,Z/_N1}  Constraint: X neq Z, dom(_N5,_N1), un(_N4,_N5,B), ran(_N5,_N3),              subset(_N3,{Y}), ...We have deliberately cut out the list of constraints returned by $\{log\}$ to simplify the exposition. pfun({[X,Y],[Z,Y]/B}) restricts the set to be a function, thus forcing X neq Z. Then, the names returned by remind includes X and Z, plus some set _N1. The constraints state that _N1 is the domain of _N5, a subset of B, whose range contains only Y. That is, Names contains the right names and nothing else. Another scenario that can be analyzed is given by Birthday being a constant function:
pfun({[X,Y],[Z,Y]/B}) & ran(B,R) & subset(R,{Y}) &remind({[X,Y],[Z,Y]/B},W,Names).Note that in this case W, not Y, is the third argument passed in to remind. We analyze only two solutions. The first one is the following:
@@ 5 Machine consistency
One of the advantages of having a state machine language in $\{log\}$ is that a collection of standard VC can be automatically generated. Once discharged, these VC ensures the specification is consistent. Then, we have implemented a Verification Condition Generator (VCG) as a $\{log\}$ component. In this section we show how the VCG works.
Once users are done with executing the prototype, they can call the VCG which checks several well-formedness conditions about the structure of the specification and in that case generates a file containing $\{log\}$ code encoding the VC as well as code to discharge them. Some of the VC generated by the VCG are the following—the examples correspond to the birthday book specification.
@@
    Invariance lemmas: each operation preservers each and every invariant. For example888addBirthday_pi_birthdayFun stands for addBirthday preserves invariant birthdayFun.:
    addBirthday_pi_birthdayFun :-  neg(    birthdayFun(Birthday) &          addBirthday(Known,Birthday,Name,Date,Known_,Birthday_)          implies    birthdayFun(Birthday_)    ).
More VC are generated when the specification uses axioms and user-defined theorems—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 13.6). The most important VC are the invariance lemmas. However, if operations or invariants are unsatisfiable, then invariance lemmas will trivially hold, thus, the VCG also generates the first two VC. Note that the first two classes of VC are satisfiability proofs whereas the third one entails unsatisfiability proofs. In general $\{log\}$ will be very fast at discharging the first two classes of VC and slower when discharging the third one.
@@ 5.1 Discharging verification conditions
Here we analyze the second situation, whereas the third one is analyzed in Section 5.2. $\{log\}$ can timeout when solving a VC simply because many decision procedures for set theory are NP-complete, although in practice they work in many cases [^9] In $\{log\}$ most decision procedures are based on set unification which has been proved to be NP-complete (Dovier et al.,, 2006).. When a VC times out, we do not know whether the VC is indeed a theorem of the specification or not. Then, we should help $\{log\}$ to assist us to find out whether the VC is provable or not. To this end, users can rerun the VC by calling check_vcs with a couple of arguments. These arguments are passed in to $\{log\}$ making it to slightly change the solving algorithm as follows.
@@
  • Execution options. This is the more promising course of action as execution options influence the constraint solving algorithm. Influence means that the algorithm can take shorter to solve a formula but it also can take longer. It is hard to predict what execution options will have a positive impact in solving a particular formula. However there are some guidelines that in general produce good results. Here we will show a few execution options—see $\{log\}$ user’s manual for more details (Rossi and Cristiá,, 2025, Section 11.1).
    • subset_unify. Implements set equality as a double set inclusion instead of implementing it by exploiting set unification (Dovier et al.,, 2006). Note that set equality is pervasive in $\{log\}$ formulas even if the formula does not use it explicitly.
@@
    • noirules. By default $\{log\}$ applies some (optional) inference rules while the goal is processed. noirules deactivates the application of these inference rules.
@@ 5.2 Analyzing undischarged verification conditions
Every time the status of a VC is Error, $\{log\}$ saves a counterexample contradicting the VC. These counterexamples can be very helpful in finding out why the proof failed. There are two kinds of counterexamples: abstract, which may include free variables; and ground, which do not contain free variables. For example, when $\{log\}$ attempts to discharge the VC of the birthday book specification, the VC named addBirthday_pi_birthdayFun fails. We can see the ground counterexample as follows101010vcgce stands for verification condition ground counter example.:
@@
[…]
@@
We have applied this methodology, for example, to analyze Chinese wall security properties (Capozucca et al.,, 2024), where we have to add 12 hypotheses in 5 VC out of 35, that is, 30 VC need no human intervention; and to verify a model of a landing gear system (Cristiá and Rossi, 2024a, ), where 60% of the VC where discharged in the first run whereas the vast majority of the remaining proofs needed only one hypothesis.

### v2

@@ 2.4 Restricted quantifiers
We have already introduced restricted quantifiers (RQ) in $\{log\}$. We have seen foreach (restricted universal quantifier, RUQ) and exists (restricted existential quantifier, REQ), in the previous sections. Here we discuss them a little more deeply. First of all, RQ in $\{log\}$ are at least as expressive as set relation algebra (Cristiá and Rossi, 2024d , Sect. 4.3), in part, due to the following characteristics. First, RQ admit quantification terms rather than simply variables. For example:
@@ 2.5 Types
Rooted in Prolog, $\{log\}$ provides essentially an untyped language based on untyped FOL and untyped set theory. In this way, a set such as $\{1,a,(3,b),\{x,q\}\}$ is perfectly legal in $\{log\}$. Untyped formalisms are not bad in themselves but types help to avoid some classes of errors (Lamport and Paulson, 1999). For this reason, we recently defined a type system and implemented a typechecker as a $\{log\}$ component (Cristiá and Rossi, 2024b ).
@@
2.6 Undefinedness
In B and Z, undefinedness arises whenever a partial function is applied to an argument that does not belong to the domain of the function. In turn, undefinedness may lead to inconsistencies. In those notations undefinedness is solved by requesting proof obligations guaranteeing that the argument belongs to the domain of the partial function.
In $\{log\}$ undefinedness is less pervasive because function application becomes a predicate rather than a term, as shown in Section 2.1. In effect, if X does not belong to the domain of F, apply(F,X,Y) is simply unsatisfiable. In other words, function application cannot remain undefined: it is either satisfiable or unsatisfiable. Same considerations hold for applyTo.
Another situation for undefinedness is division by zero in integer arithmetic. In this case B and Z requires to prove that $y\neq 0$ whenever $x\mathbin{\mathtt{div}}y$ is in context. In $\{log\}$ this situation does not occur given that it only supports linear integer arithmetic which rules out expressions such as $x\mathbin{\mathtt{div}}y$. Whenever $\{log\}$ tries to solve a predicate including such a term an exception is raised.
In summary, in $\{log\}$ function application does not lead to inconsistencies.
@@ 3 A logic language for state machines
As we have explained in the introduction, we have defined a language on top of $\{log\}$ for the description of state machines. The language is inspired in the B and Z specification languages. This language constrains programmers to use $\{log\}$ in a more restricted way but increasing the chances of automatic verification.
We will introduce the main elements of the language and the functionalities described in coming sections by means of a running example. Some elements are deliberately left out of the paper for brevity. More details can be found in the $\{log\}$ user’s manual (Rossi and Cristiá, 2025, Section 13) and in 10 case studies that are publicly available. [^7] Case studies and running example: https://www.clpset.unipr.it/SETLOG/APPLICATIONS/fv.zip. These case studies include informal requirements, the $\{log\}$ specification of a state machine, Next scenarios (Section 4.1), VCG commands (Section 5) and test case generation (Section 6). These case studies make use of some features not included in the running example such as RUQ, cardinality, integer arithmetic, set relation algebra, etc.
@@
A $\{log\}$ state machine is a collection of declarations and predicates. Predicates and declarations are similar to Prolog’s. State machines transition between states defined by state variables. The first declaration is used to declare the state variables of the state machine. In the birthday book we have the following declaration:
@@
   invariant(birthdayFun).   birthdayFun(Birthday) :- pfun(Birthday).As can be seen, each invariant is preceded by an invariant declaration. The body of each invariant is a $\{log\}$ formula depending on the state variables. An invariant declaration is a declaration of intent. It remains to be proved that these predicates are indeed preserved by the operations of the specification—see Section 5. By the end of Section 5.2 we analyze the possibility of conjoining these invariants in just one declaration.
@@
The main part of a $\{log\}$ state machine specification is the definition of the operations (state transitions) of the specification. Operations depend on state variables and input and output parameters. Besides, for each state variable $V$ the predicate can also depend on V_V\_, which represents the value of $V$ in the next state [^8] V_V\_ plays the same role than $V^{\prime}$ in Z specifications. We chose to use ‘_’ instead of the prime character due to Prolog compatibility.. The first operation of the birthday book adds a name and the corresponding birth date to the system. Hence, the head of the clause is the following:
@@
   addBirthdayOk(Known,Birthday,Name,Date,Known_,Birthday_) :-    Name nin Known &    un(Known,{Name},Known_) &    un(Birthday,{[Name,Date]},Birthday_).It is easy to see that the specification is given by providing pre- and postconditions. For instance, the constraint Name nin Known is a precondition whereas un(Birthday,{[Name,Date]},Birthday_) is a postcondition stating that the value of Birthday in the next state is equal to the union of its value in the before state with {[Name,Date]}.
@@
   operation(remind).   remind(Birthday,Today,Names) :-    rres(Birthday,{Today},M) & dom(M,Names).where rres(R,B,S) is a constraint called range restriction whose interpretation is $S=\{(x,y)\in R\mid y\in B\}$999rres/3is one of the relational operators that can be defined in terms of the classic set operators and the basic operators of set relation algebra.. Note that state variable Known is not an argument simply because the operation does not need it. Here Today is (supposed to be) an input whereas Names is (supposed to be) an output. This distinction is enforced by the user when the specification is executed. remind does not transition to a new state; it just outputs information. This is specified by not including the next-state variables in the head of the clause.
@@
3.1 Model parameters, axioms and user-defined theorems
The state machine specification language also supports parameters, axioms and user-defined theorems (Rossi and Cristiá, 2025, Section 13.1.1). Parameters play the role of machine parameters and constants in B specifications and the role of variables declared in axiomatic definitions in Z. That is, parameters serve to declare the existence of some (global) values accessible to invariants and operations, but they cannot be changed by operations—i.e., there is no next-state value for a parameter. Axioms are used to state properties of parameters. User-defined theorems are used to state properties that can be deduced from axioms, invariants, operations or theorems that have already been declared.
@@ 3.2 Typing state machines
[…]
@@ 4.1 The Next environment
Next is a component tightly integrated with $\{log\}$ that simplifies the execution and analysis of deterministic, fully instantiated functional scenarios. Although Next is less general than the approach shown in the next section, it considerably eases the job of users. For example, we can add a person’s name and birth date when the system is in its initial state, as follows:
   {log}=> initial >> addBirthday(Name:alice,Date:may24).    Known = {alice}, Birthday = {[alice,may24]}initial refers to the predicate declared as such in the specification; >> (read then) imposes a sequential order in the execution. $\{log\}$ automatically fetches state variables when calling an operation; users need to indicate values only for the input variables (e.g., Name:alice).
With Next it is easy to add more people to the system:
   {log}=> initial >>    addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]).    Known = {alice,bob}, Birthday = {[alice,may24],[bob,nov05]}In fact the above call is equivalent to the following one:
   {log}=> initial >>    addBirthday(Name:alice,Date:may24) >>    addBirthday(Name:bob,Date:nov05).As can be seen, Next automatically chains the after-state of an operation with the before-state of the next one thus relieving users from having to introduce new variables to get each state of the system.
More operations can be called by using >> to execute more complex scenarios. Moreover, the execution trace of the scenario can be analyzed by enclosing the initial state between square brackets, as follows:
   {log}=> [initial] >>    addBirthday(Name:[[alice,bob]],Date:[[may24,nov05]]) >>    remind(Today:may24,Cards).in which case the full state trace is shown as follows:
   Known = {},   Birthday = {}    ----> addBirthday(Name:alice,Date:may24)   Known = {alice}, Birthday = {[alice,may24]}    ----> addBirthday(Name:bob,Date:nov05)   Known = {alice,bob}, Birthday = {[alice,may24],[bob,nov05]}    ----> remind(Today:may24,Cards)   Known = {alice,bob}, Birthday = {[alice,may24],[bob,nov05]},   Cards = {alice}The execution starts from the initial state (Known = {}, Birthday = {}), then Alice’s birthday is added thus arriving to the state given by Known = {alice}, Birthday = {[alice,may24]}. Afterwards, Bob’s birthday is added to the book and finally we call for the set of people whose birthday is May, 24. In this case the answer is Cards = {alice} and we can also check that remind does not change the state of the system.
Whenever Next cannot fully determine the next state or an output value, an error message is issued.
   {log}=> initial >> addBirthday(Name,Date:may24).   ERROR: (next) next-state/output not sufficiently instantiated after:    addBirthday(Name,Date:may24)Since there is no value bound to Name the next state of the system remains underspecified making Next to issue an error. This helps users to analyze the behavior of the system in situations similar to the real usage.
@@ 4.2 Symbolic executions
$\{log\}$ can perform symbolic executions too (King, 1976). That is, we can execute the prototype by providing some variables instead of constants as inputs—or a mixture of both. In this way we can draw more general conclusions about the specification. In this case, though, we cannot use Next because it requires variables involved in the scenario to be completely instantiated. Actually, in symbolic executions we need to analyze the constraints returned by $\{log\}$. The following scenario analyses how remind responds to a call where Birthday has at least two equal dates in its range:
   {log}=> pfun({[X,Y],[Z,Y]/B}) & remind({[X,Y],[Z,Y]/B},Y,Names).    Names = {X,Z/_N1}    Constraint: X neq Z, dom(_N5,_N1), un(_N4,_N5,B), ran(_N5,_N3),    subset(_N3,{Y}), ...We have deliberately cut out the list of constraints returned by $\{log\}$ to simplify the exposition. pfun({[X,Y],[Z,Y]/B}) restricts the set to be a function, thus forcing X neq Z. Then, the names returned by remind include X and Z, plus some set _N1. The constraints state that _N1 is the domain of _N5, a subset of B, whose range contains only Y. That is, Names contains the right names and nothing else.
Clearly, this kind of answers is harder to analyze than the answers given by Next although the former yield more general results. To help during the symbolic analysis of the system users can combine it with ground solutions. For example, the answer to the above scenario when executed in groundsol mode (Section 2) is the following:
   X = n0, Y = n2, Z = n1, B = {}, Names = {n0,n1}We can see that Names contains just the values of X and Z, thus coinciding with the above analysis.
Another scenario that can be analyzed is given by Birthday being a constant function:
   {log}=> pfun({[X,Y],[Z,Y]/B}) & ran(B,R) & subset(R,{Y}) &    remind({[X,Y],[Z,Y]/B},W,Names).Note that in this case W, not Y, is the second argument passed in to remind. We will analyze only two solutions: (i) Obtaining the reminders for date Y, which ensures that at least X and Z are contained in the list of names; and (ii) Obtaining reminders for any other date, which will result in an empty set of names. For the first solution $\{log\}$ returns the following:
@@ 5 Machine consistency
One of the advantages of having a state machine language in $\{log\}$ is that a collection of standard VC can be automatically generated. Once discharged, these VC ensure the specification is consistent. Then, we have implemented a Verification Condition Generator (VCG) as a $\{log\}$ component. In this section we show how the VCG works.
The VCG checks several well-formedness conditions about the structure of the specification and in that case generates a file containing $\{log\}$ code encoding the VC as well as code to discharge them. Some of the VC generated by the VCG are the following—the examples correspond to the birthday book specification.
@@
    Invariance lemmas: each operation preserves each and every invariant. For example101010addBirthday_pi_birthdayFunstands for addBirthday preserves invariant birthdayFun.:
       addBirthday_pi_birthdayFun :-    neg( birthdayFun(Birthday) &    addBirthday(Known,Birthday,Name,Date,Known_,Birthday_)    implies birthdayFun(Birthday_) ).As we discuss below, some VC may be ‘weak’ in the sense that it may be necessary to add other invariants in the premise. addBirthday_pi_birthdayFun is weak.
More VC are generated when the specification uses axioms and user-defined theorems—see $\{log\}$ user’s manual for more details (Rossi and Cristiá, 2025, Section 13.6). The most important VC are the invariance lemmas. However, if operations or invariants are unsatisfiable, then invariance lemmas will trivially hold, thus, the VCG also generates the first two VC. Note that the first two classes of VC are satisfiability proofs whereas the third one entails unsatisfiability proofs. In general $\{log\}$ will discharge satisfiable VC (i.e., the first two classes) in polynomial time and it will take an exponential time when discharging unsatisfiable VC (i.e., the third class), although it works well in practice as experiments and case studies suggest (Cristiá and Rossi, 2021a ; Cristiá and Rossi, 2021c ; Cristiá and Rossi, 2024a ; Cristiá et al., 2023; Capozucca et al., 2024).
@@ 5.1 Discharging verification conditions
Here we analyze the second situation, whereas the third one is analyzed in Section 5.2. $\{log\}$ can timeout when solving a VC simply because many decision procedures for set theory are NP-complete, although in practice they work in many cases [^11] In $\{log\}$ most decision procedures are based on set unification which has been proved to be NP-complete (Dovier et al., 2006).. When a VC times out, we do not know whether the VC is indeed a theorem of the specification or not. Then, we should help $\{log\}$ to assist us to find out whether the VC is provable or not. To this end, users can rerun the VC by calling check_vcs with a couple of arguments. These arguments are passed in to $\{log\}$ making it to slightly change the solving algorithm as follows.
@@
  • Execution options. This is the more promising course of action as execution options influence the constraint solving algorithm. Influence means that the algorithm can take shorter to solve a formula but it also can take longer. It is hard to predict what execution options will have a positive impact in solving a particular formula. However there are some guidelines that in general produce good results. Here we will show a few execution options—see $\{log\}$ user’s manual for more details (Rossi and Cristiá, 2025, Section 11.1).
    • subset_unify. Implements set equality as a double set inclusion instead of implementing it by exploiting set unification (Dovier et al., 2006). Note that set equality is pervasive in $\{log\}$ formulas even if the formula does not use it explicitly.
@@
    • noirules. By default $\{log\}$ applies inference rules while the goal is processed. noirules deactivates these inference rules because in some cases their application may slow down the constraint solving process.
@@ 5.2 Analyzing undischarged verification conditions
Every time the status of a VC is Error, $\{log\}$ saves a counterexample contradicting the VC. These counterexamples can be very helpful in finding out why the proof failed. There are two kinds of counterexamples: abstract, which may include free variables; and ground, which do not contain free variables. For example, when $\{log\}$ attempts to discharge the VC of the birthday book specification, the VC named addBirthday_pi_birthdayFun fails. We can see the ground counterexample as follows121212vcgcestands for verification condition ground counter example.:
@@
This example leads us to explain an important design decision concerning the way the VCG generates invariance lemmas. Let $\{I_{k}\}_{k\in A}$ be the invariants of some specification and let $O$ be an operation. The invariance lemma for $O$ and $I_{k_{0}}$ (for some $k_{0}\in A$) is [^13] If the specification includes axioms the statement of the invariance lemma is a little bit more complex.: $(\bigwedge_{k\in A}I_{k})\land O\implies I_{k_{0}}^{\prime}$, where $I_{k_{0}}^{\prime}\triangleq I_{k_{0}}[\forall v\in\mathit{StateVars}:v^{\prime}/v]$. However, the VCG generates the following formula: $I_{k_{0}}\land O\implies I_{k_{0}}^{\prime}$. Clearly, the invariance lemmas generated by the VCG may lack some other invariants as hypotheses. This is a design decision based on the fact that unnecessary hypotheses may severely reduce the performance of an automated theorem prover. Indeed, $\{log\}$ might start rewriting some useless hypotheses rather than the useful ones, thus spending time in dead-end proof paths. Hence, we decided that users can use counterexamples to add hypotheses one by one, if necessary. Furthermore, $\{log\}$ provides the findh command family that automates the search for hypotheses. findh uses the abstract counterexamples saved by the VCG to find invariants that contradict them. These invariants are good candidates to become hypotheses as their presence surely will not generate the same counterexamples. As with execution options, analyzing counterexamples is far less involved than attempting interactive proofs. It is important to observe that if a VC times out it leaves users without any clue about the reason for that. Instead, a failed proof accompanied by a counterexample let users to work it out.
@@
We have applied this methodology, for example, to analyze Chinese wall security properties (Capozucca et al., 2024), where we have to add 12 hypotheses in 5 VC out of 35, that is, 30 VC need no human intervention; and to verify a model of a landing gear system (Cristiá and Rossi, 2024a ), where 60% of the VC where discharged in the first run whereas the vast majority of the remaining proofs needed only one hypothesis.